OpenStudy (anonymous):

Solve the quadratic equation 3x2 + 5x + 2 = 0 by factoring

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]\[(x-x_1)(x-x_2)=0\]

OpenStudy (dls):

3x^2+5x+2 3x^2+3x+2x+2 3x(x+1)+2(x+1) (3x+2)(x+1)

OpenStudy (dls):

@UnkleRhaukus "By factorizing"

OpenStudy (unklerhaukus):

hmm, i guess i did interpret the question the wrong way